OpenStudy (anonymous):

Prove that d/dx lnx = 1/x WITHOUT assuming it or using the fact d/dx e^x = e^x @amistre64 (I used to know the good proof but I forgot)

OpenStudy (amistre64):

ln(x+h) - ln(x) ------------- h

OpenStudy (anonymous):

-.-

OpenStudy (amistre64):

or is the other fashion better for this: lnb - lna ------- b-a

OpenStudy (zarkon):

what is your definition of \(\ln(x)\)

OpenStudy (anonymous):

Use the proof that uses the limit definition of E.

OpenStudy (amistre64):

limit defintion of epsilon deltas?

OpenStudy (anonymous):

\[e = \lim_{n \rightarrow \infty}(1+\frac{ 1 }{ n })^{n}\]

OpenStudy (amistre64):

i dont think i know of a proof that would use that off hand.

OpenStudy (anonymous):

Oh another question, how come e is 2.718 when the limit definition just brings a 1 as the answer?

OpenStudy (amistre64):

squeeze thrm plays a part i believe

OpenStudy (anonymous):

Because of n is infinity then 1/n is 0 and 1^n is 1for all n

OpenStudy (zarkon):

you don't get to dictate how fast each individual n goes to infinity....they all go to infinity at the same rate

OpenStudy (amistre64):

yeah ... that logic :) ((n+1)/n)^n (2/1)^1 = 2 (3/2)^2 = 2.25 (4/3)^3 = 2.370370... (5/4)^4 = 2.441406... etc (101/100)^100 = 2.7048... the limit of the sequence evens out to e